# Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning

### v4

@@ Abstract
Metamaterials with functional responses can exhibit varying properties or functionalities under different conditions (e.g., wave-based responses or deformation-induced property variation). This work addresses the rapid inverse design of such metamaterials to meet target qualitative functional behaviors. This inverse problem is challenging due to its intractability and non-unique solutions. Past works mainly focus on deep-learning-based methods that are data-demanding, require time-consuming training and hyperparameter tuning, and are non-interpretable. We propose the Random-forest-based Interpretable Generative Inverse Design (RIGID), an iteration-free, single-shot inverse design method for fast generation of metamaterial designs with on-demand functional behaviors. RIGID leverages the interpretability of random forests, eliminating the need to train a response-design inverse model. Based on the likelihood of target satisfaction derived from the trained design-response forward model, one can sample a desired number of design solutions using Markov chain Monte Carlo methods. We validate RIGID on acoustic and optical metamaterial design problems where only small datasets (fewer than 250 training samples) are available. Synthetic design problems are created to illustrate the mechanism of likelihood estimation. This work offers a new perspective on solving on-demand inverse design problems, showcasing the potential for incorporating interpretable machine learning into generative design and eliminating its large data requirement.
@@ 1 Introduction
Functional responses refer to varying properties or behaviors under different conditions. One example is metamaterials whose electromagnetic, acoustic, or elastic wave propagation behaviors change with wavelengths or frequencies [9]. Another example is metamaterials that exhibit changing properties or functionalities due to deformation in response to external stimuli like temperature [10] and magnetic fields [11]. Tailoring functional responses of these metamaterials is of interest to applications such as sound and vibration control, analog computing, medical imaging, sensing, communication, and soft robotics. In many use cases, rather than precisely controlling the complete functional responses, we only care about qualitative behaviors under certain conditions. For example, acoustic metamaterials were usually designed to have bandgaps at specified frequencies to achieve functionalities like wave-guiding [12, 13], focusing [14, 15], and vibration mitigation [16, 17, 18]. However, it is unnecessary and computationally expensive to design for the whole dispersion relation [19, 20, 21, 22, 23]. Similarly, we may design optical metamaterials to qualitatively manipulate optical properties (e.g., high or low absorption/reflection/transmission) under certain wavelengths, without requiring the entire spectral response to match an exact target [24, 25].
@@
Identifying metamaterial designs from a given target forms an inverse design problem. Unlike many forward problems where one can obtain solutions (e.g., spectral responses or material properties under external stimuli) by modeling the physics or conducting experiments, inverse design problems are usually intractable. Traditionally, these problems are solved by iterative optimization (i.e., minimizing the difference between the actual quantity of interest and the target) [19, 20, 26, 22, 27]. This, however, requires repeatedly updating the design solution and solving forward problems. When the design target changes, one needs to rerun the entire optimization process. Thus, inverse design by iterative optimization becomes impractical if solving the forward problem (by simulations or experiments) is time-consuming or if the design target needs to change frequently. For example, in the case of acoustic metamaterial design, one may want to obtain a material waveguide containing spatially-varying unit cells with different bandgap ranges at different locations in a tessellation. To accelerate the optimization approach, prior works replaced simulations or experiments with machine learning models [28, 29]. However, the efficiency and quality of final solutions are highly dependent on both the machine learning model and the optimization algorithm. On the other hand, a single run of optimization usually only returns one final solution, although multiple designs might satisfy a given target (i.e., the non-uniqueness of solutions). For example, multiple acoustic metamaterial designs may have bandgaps within the same target frequency range. This non-uniqueness nature of inverse design problems was also shown for optical metasurfaces [30, 31, 32, 27], especially when we only care about qualitative target responses. Optimization approaches cannot explore diverse alternative solutions efficiently.
@@ 2 Methods
The next step is to trace up the tree from the $N$ relevant leaves, obtained by Step 2, to the root node (Fig. 2, Step 3). This will result in $N$ decision paths, along which are nodes indicating splitting criteria for design variables $\mathbf{x}$. Thus, each decision path represents a set of design variable ranges, or in other words, a region in the design space. We assign each region a score equal to the predicted probability at each corresponding leaf. This probability is learned from the training data and equals the proportion of positive data in a leaf. It indicates the tree’s belief in the probability of a design $\mathbf{x}$ satisfying the target $\mathcal{T}=\{f(\mathbf{x}^{*},s)=1|\forall s\in\Omega\}$ ($\Omega=[3,4]\cup[6,7]$ in this example) if the design falls in the design space region corresponding to the leaf. Therefore, with a single decision tree $i$, we already have the map of likelihood $\mathcal{L}_{m}(\mathbf{x}|\mathcal{T})=\mathbb{P}_{m}(\mathcal{T}|\mathbf{x})$ for target satisfaction: each of the $N$ regions has a uniformly distributed likelihood equal to the predicted probability at the corresponding leaf, and the rest of the design space has a likelihood of 0 (Fig. 2, Step 3).
@@
In this work, we assume a uniform prior, which is equivalent to just considering the likelihood $\mathcal{L}(\mathbf{x}|\mathcal{T})$.
@@ 3 Results
We compare RIGID to a commonly used low-cost generative design strategy—surrogate modeling combined with the genetic algorithm (GA) [63, 64, 65, 66, 67]. We choose this baseline according to the following criteria to ensure that its efficiency and application scenario are comparable to those of RIGID: 1) the baseline method should generate design solutions quickly (i.e., no additional training or design evaluation is required during the use phase), 2) the baseline method can generate multiple satisfying solutions to the inverse design problem, and 3) the baseline method can handle small data. These criteria exclude sequential design strategies like Bayesian optimization (as they violate the first criterion) and any existing single-shot inverse design methods mentioned in Sec. 1 (as direct inverse mapping and T-NN violate the second criterion and conditional generative model-based methods usually violate the last criterion).
The genetic algorithm is an optimization method that iteratively evolves a population of candidate solutions, based on biologically inspired operators such as mutation, crossover, and selection, toward better solutions [68]. As we usually need to evaluate a large population of candidate solutions at each iteration (referred to as a generation), a surrogate model (usually a machine learning model) is often constructed to accelerate design response prediction, thereby speeding up the optimization process. For the surrogate model in the baseline method, we use the same trained random forest as in RIGID for consistency, eliminating differences due to forward prediction accuracy and focusing on comparing the inverse design methods. We set the objective of GA to maximize the aggregated probability of target satisfaction within the target domain $\Omega$, i.e., $\max_{\mathbf{x}}\sum_{s_{i}\in\Omega}f^{\prime}(\mathbf{x},s_{i})$, where $\{s_{i}\}_{1}^{n}$ are discretized auxiliary variable values (e.g., discrete values of frequencies or wavelengths) and $f^{\prime}$ represents the trained random forest. The initial population for GA is generated by randomly sampling solutions within the design variable bounds. At each generation, we select individuals to reproduce for a new generation by randomly picking three individuals from the population, selecting the best one, and repeating this process 300 times. We set the probability for crossover and mutation to be 0.5 and 0.2, respectively. Once an individual is mutating, there is a 5% chance for each design variable in this individual to mutate. For a continuous design variable, we define the mutation operation as adding Gaussian noise with a mean of 0 and a standard deviation equal to 1% of the variable’s range. For a categorical design variable, mutation is defined as randomly reassigning the category of the variable. We ran GA for 10 generations, each with a population of 300. Finally, we select a set of solutions with the highest objective values among all the individuals created during GA, and compare them to designs generated by RIGID. From this point, we will call designs generated by RIGID and GA as RIGID designs and GA designs, respectively.
@@ 3.1 Applying RIGID to Design Acoustic Metamaterials with Target Bandgaps
Figure 3: Acoustic metamaterial design problem configuration and results. (A) Design variables of center and corner mass radii ($r_{\text{center}}$ and $r_{\text{corner}}$) and strut radius ($r_{\text{strut}}$). (B) High symmetry points of the cubic irreducible Brillouin zone. (C) A sample dispersion relation and bandgap (marked by the highlighted zone). The design objective is to generate new acoustic metamaterial designs with target bandgaps. (D) KDE of the estimated likelihood for generated designs. (E) Satisfaction rates, average scores, and selection rates for RIGID designs under varying sampling thresholds (solid lines), in comparison to the satisfaction rate of GA designs (horizontal dashed line). The horizontal dotted line indicates 100% satisfaction (F) Geometries and corresponding dispersion relations of five RIGID designs with the highest likelihood of satisfying a specified target bandgap (marked as highlighted frequency regions). All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m. Here only the fourth design fails to meet a small portion (at around 6 MHz) of the target bandgap, whereas the others meet it. Generated designs for other targets are shown in SI Appendix, Figs. S2-S6.
@@
Figure 4: Distributions of satisfactory solutions for two bandgap targets. The left panel shows that GA designs are highly localized while RIGID can lead to diverse solutions. The right panel indicates that none of the GA designs satisfy the target, while satisfactory RIGID designs are diverse and can be very different from feasible designs from data. Solutions from data include feasible designs in both training and test data.
@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
We also generated the same amount of designs under the same targets by using GA. We found that GA designs can generate 100% satisfaction rate (Fig. 5E). However, similar to the acoustic metamaterial design example, the generated designs are localized compared to RIGID designs which cover much wider feasible region(s) in the design space. This can be reflected by the design variable distributions of generated feasible solutions shown in Fig. 5G, which corresponds to generated designs for a randomly created target (the target is shown as the highlighted wavelength regions in Fig. 5F). For example, all the GA designs have the same cross-section geometric type while RIGID generated satisfactory designs with different geometric types. It is easy to see that, without considering solution diversity, achieving a high target satisfaction rate is trivial—as long as we discover one satisfactory solution, we can add small perturbations to that solution to generate an infinite number of alternative solutions that have similar responses and thus are also likely to satisfy the target. Note that compared to the acoustic problem, this optical problem is less challenging as there is an average of 28.5% data already satisfying the ten targets (in contrast to an average of 9.3% satisfactory designs from data in the acoustic problem). Therefore, finding a few satisfactory solutions for the optical problem is relatively easy; the challenge lies in discovering all feasible solutions, where RIGID excels by identifying a wide range of them. We can also see that when the sampling threshold is 0.8, RIGID achieves a satisfaction rate of 100%, demonstrating its ability to trade the diversity of generated designs for higher feasibility.
@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
We also visualize the distribution of satisfactory designs generated by both RIGID and GA in Fig. 7A. Again, we observe a much broader coverage of the feasible design space by RIGID designs, while GA designs are merely minor perturbations of one or two feasible solutions.
@@ 4.1 Interpretability
RIGID offers two levels of interpretability. Firstly, RIGID is more transparent than deep learning models. The transparency of the random forest as a forward model helps us understand how the predictions are made. The inverse design process based on likelihood estimation and sampling is also transparent. The transparency of both forward and inverse processes offers insights into the decision-making process and diagnostic capabilities. Secondly, owing to the interpretability of the forward prediction model, we can derive the likelihood, which the inverse design relies on, by probing into each decision path of trees in the random forest. The likelihood is an interpretable metric to help us understand the probability of a generated design satisfying the given target and thus aid the decision-making process when selecting final solutions. The results (Figures 3D, 3E, 5D, 5E, and 7C) have demonstrated a high correlation between the likelihood and actual probability of target satisfaction. The likelihood also serves as an indicator of the need for adjusting the design space. When the likelihood is low across the entire design space, the model believes the target is almost impossible to achieve, which may indicate a need for expanding the design space (i.e., extending design space bounds or adding new design freedom) to allow wider coverage of the response space and easier discovery of satisfying solutions. Please see more discussion in SI Appendix, Sec. S4. Another benefit of having the explicit likelihood as an interpretable metric is that we can adjust the likelihood threshold to trade-off exploration and exploitation — higher (lower) threshold means more focus on exploitation (exploration). Both levels of interpretability were absent from most prior works of iteration-free, single-shot inverse design methods for metamaterials.
@@ 4.2 Data Demand
RIGID is particularly useful when data collection is expensive, as in many cases where high-fidelity simulation or experimental data are needed. Although hardware-accelerated simulation and surrogate solvers have advanced in recent years to make simulation faster without sacrificing too much accuracy, the trade-off between time/computational resources and accuracy will always pose a problem. For example, hardware acceleration for high-fidelity simulation relies on high-end GPUs/CPUs and other high-performance computing infrastructure, which might not be accessible under limited budgets. With surrogate solvers, the need for computational resources reduces, but so will the accuracy. If the surrogate solver is data-driven, we will still have the problem of collecting high-fidelity data in the first place. Thus, we believe that the issue of data deficiency will always exist, especially when high-fidelity data is desired but time and computational resources are constrained. With most of the prior inverse design works based on deep learning with a high demand for data and computational resources, it is useful to create a method that works under scenarios at the other end of the data/computational resource requirement spectrum. Additionally, our proposed method is not only useful with high-fidelity simulation/numerical data, but also with experimental data. When it is difficult to build an accurate numerical model, the only way to acquire high-fidelity data is through experiments. In this case, data deficiency will become a pressing challenge for deep learning-based methods.
This work focuses on the scenario of small data, low-dimensional problems. As the problem dimension increases, the requirement for dataset size will quickly increase due to the curse of dimensionality. While there is a large collection of deep learning methods to address high-dimensional and large dataset problems, the setting of small data and low dimensionality seems to be under-studied by prior works, despite this setting’s practicality under constrained budgets. RIGID fills this gap by offering a much more data-efficient and interpretable alternative to deep-learning-based single-shot inverse design methods.

### v6

@@ Abstract
[…]
@@ 1 Introduction
Functional responses refer to varying properties or behaviors under different conditions. A notable example that involves functional responses is metamaterials, which alter their electromagnetic, acoustic, or elastic wave propagation behaviors depending on the wavelength or frequency [9]. Another example is metamaterials that exhibit changing properties or functionalities due to deformation in response to external stimuli like temperature [10] and magnetic fields [11]. Tailoring functional responses of these metamaterials is of interest to applications such as sound and vibration control, analog computing, medical imaging, sensing, communication, and soft robotics. In many use cases, rather than precisely controlling the complete functional responses, we only care about qualitative behaviors under certain conditions. For example, acoustic metamaterials were usually designed to have bandgaps at specified frequencies to achieve functionalities like wave-guiding [12, 13], focusing [14, 15], and vibration mitigation [16, 17, 18]. However, it is unnecessary and computationally expensive to design for the whole dispersion relation [19, 20, 21, 22, 23]. Similarly, we may design optical metamaterials to qualitatively manipulate optical properties (e.g., high or low absorption/reflection/transmission) under certain wavelengths, without requiring the entire spectral response to match an exact target [24, 25].
@@
Identifying metamaterial designs from a given target forms an inverse design problem. Unlike many forward problems where one can obtain solutions (e.g., spectral responses or material properties under external stimuli) by modeling the physics or conducting experiments, inverse design problems are usually intractable. Traditionally, these problems are solved by iterative optimization (i.e., minimizing the difference between the actual quantity of interest and the target) [19, 20, 26, 22, 27]. This, however, requires repeatedly updating the design solution and solving forward problems. When the design target changes, one needs to rerun the entire optimization process. Thus, inverse design by iterative optimization becomes impractical if solving the forward problem (by simulations or experiments) is time-consuming or if the design target needs to change frequently. For example, in the case of acoustic metamaterial design, one may want to obtain a material waveguide containing spatially-varying unit cells with different bandgap ranges at different locations in a tessellation. To accelerate the optimization approach, prior works replaced simulations or experiments with machine learning models [28, 29]. However, the efficiency and quality of final solutions are highly dependent on both the machine learning model and the optimization algorithm. On the other hand, a single run of optimization usually only returns one final solution, although multiple designs might satisfy a given target (i.e., the non-uniqueness of solutions). For example, multiple acoustic metamaterial designs may have bandgaps within the same target frequency range. This non-uniqueness nature of inverse design problems was also shown for optical metasurfaces [30, 31, 32, 27], especially when we only care about qualitative target responses. It is generally advantageous for inverse design approaches to yield a diverse set of solutions, as this allows consideration of additional figures of merit beyond target satisfaction. For example, deriving diverse solutions for acoustic metamaterials will allow us to consider the kinematic compatibility of multiple designs in waveguide design problems. However, optimization approaches cannot explore diverse alternative solutions efficiently.
@@ 2 Methods
The next step is to trace up the tree from the $N$ relevant leaves, obtained by Step 2, to the root node (Fig. 2, Step 3). This will result in $N$ decision paths, along which are nodes indicating splitting criteria for design variables $\mathbf{x}$. Thus, each decision path represents a set of design variable ranges, or in other words, a region in the design space. We assign each region a score equal to the predicted probability at each corresponding leaf. This probability is learned from the training data and equals the proportion of positive data in a leaf. It indicates the tree’s belief in the probability of a design $\mathbf{x}$ satisfying the target $\mathcal{T}=\{f(\mathbf{x}^{*},s)=1|\forall s\in\Omega\}$ ($\Omega=[3,4]\cup[6,7]$ in this example) if the design falls in the design space region corresponding to the leaf. Therefore, with a single decision tree $m$, we already have the map of likelihood $\mathcal{L}_{m}(\mathbf{x}|\mathcal{T})=\mathbb{P}_{m}(\mathcal{T}|\mathbf{x})$ for target satisfaction: each of the $N$ regions has a uniformly distributed likelihood equal to the predicted probability at the corresponding leaf, and the rest of the design space has a likelihood of 0 (Fig. 2, Step 3).
@@
In this work, we assume a uniform prior, making the posterior directly proportional to the likelihood $\mathcal{L}(\mathbf{x}|\mathcal{T})$, and sampling from the posterior is equivalent to sampling from the likelihood.
@@ 3 Results
We compare RIGID to a commonly used low-cost generative design strategy—surrogate modeling combined with a genetic algorithm (GA) [63, 64, 65, 66, 67]. We choose this baseline according to the following criteria to ensure that its efficiency and application scenario are comparable to those of RIGID: 1) the baseline method should generate design solutions quickly (i.e., no additional training or design evaluation is required given different design targets), 2) the baseline method can generate multiple satisfying solutions to the inverse design problem, and 3) the baseline method can handle small data. These criteria exclude sequential design strategies like Bayesian optimization (as they violate the first criterion) and any existing single-shot inverse design methods mentioned in Sec. 1 (as direct inverse mapping and T-NN violate the second criterion and conditional generative model-based methods usually violate the last criterion).
Genetic algorithms are optimization methods that iteratively evolve a population of candidate solutions, based on biologically inspired operators such as mutation, crossover, and selection, toward better solutions [68]. As we usually need to evaluate a large population of candidate solutions at each iteration (referred to as a generation), a surrogate model (usually a machine learning model) is often constructed to accelerate design response prediction, thereby speeding up the optimization process. For the surrogate model in the baseline method, we use the same trained random forest as in RIGID for consistency, eliminating differences due to forward prediction accuracy and focusing on comparing the inverse design methods. We set the objective of GA to maximize the aggregated probability of target satisfaction within the target domain $\Omega$, i.e., $\max_{\mathbf{x}}\sum_{s_{i}\in\Omega}f^{\prime}(\mathbf{x},s_{i})$, where $\{s_{i}\}_{i=1}^{n}$ are discretized auxiliary variable values (e.g., discrete values of frequencies or wavelengths) and $f^{\prime}$ represents the trained random forest. We include the detailed configurations of GA in Sec. 6. After optimization, we rank the solutions generated throughout the optimization process based on their objective values and select the top solutions as final generated designs. We compare these designs to those generated by RIGID. From this point, we will call designs generated by RIGID and GA as RIGID designs and GA designs, respectively.
@@ 3.1 Applying RIGID to Design Acoustic Metamaterials with Target Bandgaps
Figure 3: Acoustic metamaterial design problem configuration and results. (A) Design variables of center and corner mass radii ($r_{\text{center}}$ and $r_{\text{corner}}$) and strut radius ($r_{\text{strut}}$). (B) High symmetry points of the cubic irreducible Brillouin zone. (C) A sample dispersion relation and bandgap (marked by the highlighted zone). The design objective is to generate new acoustic metamaterial designs with target bandgaps. (D) KDE of the estimated likelihood for generated designs. (E) Satisfaction rates, average scores, and selection rates for RIGID designs under varying sampling thresholds (solid lines), in comparison to the satisfaction rate of GA designs (horizontal dashed line). The horizontal dotted line indicates 100% satisfaction. (F) Geometries and corresponding dispersion relations of five RIGID designs with the highest likelihood of satisfying a specified target bandgap (marked as highlighted frequency regions). All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m. Here only the fourth design fails to meet a small portion (at around 6 MHz) of the target bandgap, whereas the others meet it. Generated designs for other targets are shown in SI Appendix, Figs. S2-S6.
@@
Figure 4: Distributions of satisfactory solutions for two bandgap targets. The off-diagonal plots show the pairwise bivariate distributions of design variables, and the diagonal plots show the marginal distributions of the data in each column. The left panel shows that GA designs are highly localized while RIGID can lead to diverse solutions. The right panel indicates that none of the GA designs satisfy the target, while satisfactory RIGID designs are diverse and can be very different from feasible designs from data. Solutions from data include feasible designs in both training and test data.
@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
We also generated the same amount of designs under the same targets by using GA. We found that GA designs can generate 100% satisfaction rate (Fig. 5E). However, similar to the acoustic metamaterial design example, the generated designs are localized compared to RIGID designs which cover much wider feasible region(s) in the design space. This can be reflected by the design variable distributions of generated feasible solutions shown in Fig. 5G, which corresponds to generated designs for a randomly created target (the target is shown as the highlighted wavelength regions in Fig. 5F). For example, all the GA designs have the same cross-section geometric type while RIGID generated satisfactory designs with different geometric types. It is easy to see that, without considering solution diversity, achieving a high target satisfaction rate is trivial—as long as we discover one satisfactory solution, we can add sufficiently-small perturbations to that solution to generate an infinite number of alternative solutions that have similar responses and thus are also likely to satisfy the target. Note that compared to the acoustic problem, this optical problem is less challenging as there is an average of 28.5% data already satisfying the ten targets (in contrast to an average of 9.3% satisfactory designs from data in the acoustic problem). Therefore, finding a few satisfactory solutions for the optical problem is relatively easy; the challenge lies in discovering all feasible solutions, where RIGID excels by identifying a wide range of them. We can also see that when the sampling threshold is 0.8, RIGID achieves a satisfaction rate of 100%, demonstrating its ability to trade off the diversity of generated designs for higher feasibility.
@@ 3.3 Synthetic Design Problems for Rapid Validation and Visualization
We also visualize the distributions of satisfactory designs generated by both RIGID and GA in Fig. 7A. Again, we observe a much broader coverage of the feasible design space by RIGID designs, while GA designs are merely minor perturbations of one or two feasible solutions.
@@ 4.1 Interpretability
RIGID offers two levels of interpretability. Firstly, RIGID is more transparent than deep learning models. The transparency of the random forest as a forward model helps us understand how the predictions are made. The inverse design process based on likelihood estimation and sampling is also transparent. The transparency of both forward and inverse processes offers insights into the decision-making process and diagnostic capabilities. Secondly, owing to the interpretability of the forward prediction model, we can derive the likelihood, which the inverse design relies on, by probing into each decision path of trees in the random forest. The likelihood is an interpretable metric to help us understand the probability of a generated design satisfying the given target and thus aid the decision-making process when selecting final solutions. The results (Figures 3D, 3E, 5D, 5E, and 7C) have demonstrated a high correlation between the likelihood and actual probability of target satisfaction. The likelihood also serves as an indicator of the need for adjusting the design space. When the likelihood is low across the entire design space, the model believes the target is almost impossible to achieve, which may indicate a need for expanding the design space (i.e., extending design space bounds or adding new design freedom) to allow wider coverage of the response space and easier discovery of satisfying solutions. Please see more discussion in SI Appendix, Sec. S4. Another benefit of having the explicit likelihood as an interpretable metric is that we can adjust the likelihood threshold to trade-off between exploration and exploitation — higher (lower) threshold means more focus on exploitation (exploration). Both levels of interpretability were absent from most prior works of iteration-free, single-shot inverse design methods.
@@ 4.2 Data Demand
RIGID is particularly useful when data collection is expensive, as in many cases where high-fidelity simulation or experimental data are needed. Despite the advancement of surrogate solvers and hardware-accelerated simulation, a persistent trade-off between computational resources and accuracy remains, exacerbated by challenges in accessing high-performance infrastructure and collecting high-fidelity data. With most of the prior inverse design works based on deep learning with a high demand for data and computational resources, it is useful to create a method that works under scenarios at the other end of the data/computational resource requirement spectrum. Additionally, when the only way to acquire high-fidelity data is through experiments, data deficiency will become a pressing challenge for deep learning-based methods. While there is a large collection of deep learning methods to address single-shot inverse design problems with large datasets, the setting of small data seems to be under-studied by prior works, despite this setting’s practicality under constrained budgets. RIGID fills this gap by offering a much more data-efficient and interpretable alternative to deep-learning-based single-shot inverse design methods.